# Universality of almost periodicity in bounded discrete time series

### 01 Mar 2024 v4

@@ Abstract.
almost periodic function, chaos in discrete time, recurrent neural network
@@ 1. Introduction
Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series arbitrarily chosen, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}$ such that (for a technical reason, we employ transcendental numbers)
  • $\{a_{1}^{K},a_{2}^{K},\cdots,a^{K}_{K-1},a_{K}^{K}\}\subset\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\subset\mathbb{R}\setminus\mathbb{Q}$,
@@
We discretize $y$ as follows:
@@
where $a-0:=a-\varepsilon$ for any sufficiently small $\varepsilon>0$. In what follows we sometimes omit the index $K$ unless confusion occurs. To define a suitable statistical feature, we need to classify patterns in $\bar{y}$. Let $\sigma_{n}$ ($n=1,2,\cdots,N$) be a permutation operator, namely,
@@
where $\bar{u}$ is generated by (12), and let $\Phi:[-1,1]^{L}\to[-1,1]^{L}$ be such that $Y_{t-1}\mapsto Y_{t}$ (in Appendix 3, we give the simplest example of $\Phi$). We see that, for any $t\in\mathbb{Z}_{\geq 0}$, there exists $t^{*}\in\mathbb{Z}_{\geq 0}$ such that
@@ 2. Proof of main theorem
for $\ell\in\{1,\cdots,L\}$ and $n\in\{1,\cdots N\}$, where $\sigma^{*}_{i}(0):=\sigma_{i}^{*}(L)$. The definition of the following $N\times N$ matrix $X$ is the key to prove the main theorem:
@@
$X$ is a regular matrix provided by appropriately chosen $h:\mathbb{R}\to[-1,1]$.
@@
First, by the Lindemann–Weierstrass theorem, we see that
(8) $\frac{a_{k^{\prime}}}{a_{k}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-1\}\quad(k\not=k^{\prime}),$
(9) $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}}{a_{k_{\ell}}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-L,-L+1,\cdots,-1,0,1,\cdots,L-1\}$
except for the $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$ case. Also
$\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}=L$
if and only if $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$. By (7), (8), (9) and Lindemann–Weierstrass theorem, we see
$\sum_{\ell=1}^{L}\sigma^{*}_{j}(\ell-1)\sigma_{j}(\ell)=L\quad\text{and}\quad\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\not=L\quad(i\not=j).$
In order to construct an appropriate $h$, we use a finite set $G$ as follows:
$G:=\left\{\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell):i,j\in\{1,2,\cdots,N\}\right\}\subset\mathbb{R}.$
Note that $L\in G$. Now we take a smooth function $h:\mathbb{R}\to[-1,1]$ satisfying the following:
$\begin{split}&h(L)\in\mathbb{Q}\setminus\{0\},\\ &h(\gamma)\in\left(\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\cup\{0\}\right)\subset(\mathbb{R}\setminus\mathbb{Q})\cup\{0\},\quad\text{for}\quad\gamma\subset G\setminus\{1\}.\end{split}$
Then we can easily check that (applying the Lindemann–Weierstrass theorem)
  • $h(\gamma_{1})h(\gamma_{2})\cdots h(\gamma_{N})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}$  for  $\{\gamma_{n}\}_{n=1}^{N}\in G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
  • for any $\{\tau_{n^{\prime}}\}_{n^{\prime}=1}^{N!-1}\in\{-1,1\}^{N!-1}$ and $\{\{\gamma_{n,n^{\prime}}\}_{n=1}^{N}\}_{n^{\prime}=1}^{N!-1}\subset G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
    $\sum_{n^{\prime}=1}^{N!-1}\tau_{n^{\prime}}h(\gamma_{1,n^{\prime}})h(\gamma_{2,n^{\prime}})\cdots h(\gamma_{N,n^{\prime}})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
By applying the above properties, then we see that the determinant of the matrix $X$ is nonzero, since it is expressed as
$|X|=\eta_{1}+\eta_{2},\quad\eta_{1}\in\mathbb{Q}\setminus\{0\},\quad\eta_{2}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
Now we resume the proof of the main theorem. Let us define $N\times N$ matrices: $Y$ and $W$ as follows:
(10) $Y:=\left[\sum_{\ell=1}^{L}\sigma_{i}^{*}(\ell)\sigma_{j}(\ell)\right]_{i,j}\quad\text{and}\quad W:=YX^{-1}.$
@@
(12) $\begin{split}&\begin{cases}\bar{r}(t_{0})&=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&=W^{out}\bar{r}(t),\end{cases}\end{split}$
for $t=t_{0}+1,t_{0}+2,\cdots$, where $W^{out}:=AX^{-1}$ and $A:=(a_{k(1)},a_{k(2)},\cdots,a_{k(N)})$. Rigorously we regard $\bar{u}(t-L-1)$ as $\bar{y}(t-L-1)$ for $t=t_{0}+1,t_{0}+2\cdots,t_{0}+L$. Now we show that $\bar{u}_{K}(t)$ tends to a discretized periodic function for any $K$. By (3), for each $K$, we have a chain of the permutation operator $(\sigma_{j},a_{k(j)})$ as follows:
@@
Note that the sequence $j_{1},j_{2},j_{3}\cdots$ is different from $n_{1},n_{2},n_{3}\cdots$ in (4). Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite, then there exist $m,M\in\mathbb{Z}_{\geq 2}$ ($m<M$) such that $(\sigma_{j_{m}}a_{k(j_{m})})=(\sigma_{j_{M}},a_{k(j_{M})})$. This means that there exists $T_{K}$ such that $\bar{u}_{K}(t)$ is a periodic chain for $t>T_{K}$. Moreover $\lim_{K\to\infty}\bar{u}_{K}=:u$ is uniform convergence due to (4). Thus $u(t)$ tends to an almost periodic function for $t\to\infty$ (see Corduneanu [3, (1.2)]).
Remark 2.
In some situations, this periodic chain may be different for different choice of $t_{0}$, but we do not go into the detail here.
@@ 3. Appendix
Acknowledgments. The author is grateful to Professors Masanobu Inubushi, Chikara Nakayama and Yoshitaka Saiki for valuable comments. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 20H01819.

### 25 Mar 2024 v5

@@ Abstract.
almost periodic function, chaos in discrete time, recurrent neural network, topological ring
@@ 1. Introduction
Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series arbitrarily chosen, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}\subset[-1,1]$ such that
@@
  • $a_{i}^{K}\not=0$ for any $1\leq i\leq K$ and
    $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}\not=L$ for each $(k_{1},\ldots,k_{L})\not=(k^{\prime}_{1},\ldots,k^{\prime}_{L})$,
@@
Note that the above second condition is almost always satisfied in the following sense.
Proposition 1.
There is a closed set $Z$ of $[-1,1]^{K}$ whose Lebesgue measure is zero such that the above second condition is satisfied as soon as $(a^{K}_{1},\ldots,a^{K}_{K})$ is not in $Z$.
Proof.
We drop the superscipt $K$ for simplicity. Regard $a_{i}$ ($1\leq i\leq K$) as indeterminates. Clearly the measure of the set where $a_{i}=0$ is zero. Under the assumption that any $a_{i}$ is not zero, the condition $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}}{a_{k_{\ell}}}\not=L$ is equivalent to $\sum_{\ell}a_{k_{1}}\cdots a_{k^{\prime}_{\ell}}\cdots a_{k_{L}}-L\prod_{\ell}a_{k_{\ell}}\not=0$, the left-hand-side of which is a nonzero polynomial whenever $k_{\ell}\not=k^{\prime}_{\ell}$ for some $\ell$ (then the term $\prod_{\ell}a_{k_{\ell}}$ does not vanish). Hence by the following simple fact, the inequality holds for almost all values of $a_{i}$:
$\bullet$ Let $f$ be a nonzero polynomial in $M$ variables. Then the Lebesgue measure of the zero point set $\{x\in\mathbb{R}^{M}:f(x)=0\}$ is zero.
We now discretize $y$ as follows:
@@
where $a-0:=a-\varepsilon$ for any sufficiently small $\varepsilon>0$. In what follows we sometimes omit the index $K$ unless confusion occurs (for example, $\bar{y}=\bar{y}_{K}$ and $\bar{u}=\bar{u}_{K}$). To define a suitable statistical feature, we need to classify patterns in $\bar{y}$. Let $\sigma_{n}$ ($n=1,2,\cdots,N$) be a permutation operator, namely,
@@
where $\bar{u}$ is generated by (10), and assume there exists a map $\Phi:[-1,1]^{L}\to[-1,1]^{L}$ such that $Y_{t-1}\mapsto Y_{t}$ (in Appendix 4, we give the simplest example of $\Phi$). Then we see that, for any $t\in\mathbb{Z}_{\geq 0}$, there exists $t^{*}\in\mathbb{Z}_{\geq 0}$ such that
@@ 2. Proof of main theorem
for $\ell\in\{1,\cdots,L\}$ and $n\in\{1,\cdots,N\}$, and $\sigma^{*}_{i}(L):=\sigma_{i}^{*}(0)$. The definition of the following $N\times N$ matrix $X$ and a finite set $G\subset\mathbb{R}$ are the keys to prove the main theorem:
@@
$G:=\left\{\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell):i,j\in\{1,2,\cdots,N\}\right\}.$
Note that $L\in G$.
@@
$X$ is a regular matrix for almost all $h\colon\mathbb{R}\to[-1,1]$ in the following sense. Let $\gamma_{1},\ldots,\gamma_{M}$ be the all elements of $G$. Then there is a closed set $Z$ of $\mathbb{R}^{M}$ whose Lebesgue measure is zero such that $X$ is regular as soon as $(h(\gamma_{1}),\ldots,h(\gamma_{M}))$ is not in $Z$.
@@
Take a bijection $x\colon\gamma_{m}\mapsto x_{m}$ from $G$ to the set of indeterminates $\{x_{1},\ldots,x_{M}\}$. Consider the determinant $D$ of the matrix $[x(\sum\sigma_{i}^{*}(\ell-1)\sigma_{j}(\ell))]_{i,j\in N}$ with polynomial coefficients. Then $D$ is a nonzero polynomial because it is a monic polynomial of degree $N$ with respect to the indeterminate $x(L)$ (the leading term $x(L)^{N}$ comes from the diagonal). Now we apply the simple fact in the proof of Proposition 1 to $D$. Let $Z$ be the zero point set of $D$. Assume that $(h(\gamma_{1}),\ldots,h(\gamma_{M}))$ is not in $Z$. Then $\det X$ defined by $h$ is not zero so that $X$ is regular. ∎
With the aid of the inverse of $X$, we can define a $N\times N$ matrix $W$ as follows:
(8) $W:=YX^{-1}\quad\text{with}\quad Y:=\left[\sum_{\ell=1}^{L}\sigma_{i}^{*}(\ell)\sigma_{j}(\ell)\right]_{i,j}.$
Remark 2.
There is a smooth function $h\colon\mathbb{R}\to\mathbb{R}$ such that the absolute value of any eigenvalue of the $W$ and $W^{T}W$ are strictly less than $1$ (c.f. [3]). This can be shown as follows: Let $c$ be a real number which is strictly greater than the absolute value of any eignevalue of $Y$ and the square root of the absolute value of any eigenvalue of $Y^{T}Y$. Then we can take a smooth function $h$ whose support does not contain any $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}$ for $(k_{1},\ldots,k_{L})\not=(k^{\prime}_{1},\ldots,k^{\prime}_{L})$ such that $h(L)=c$. Then $X=cI$, where $I$ is the unit matrix, and $W=\frac{1}{c}Y$, $W^{T}W=c^{-2}Y^{T}Y$.
@@
(10) $\begin{split}&\begin{cases}\bar{r}(t_{0})&:=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell-1)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&:=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&:=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&:=W^{out}\bar{r}(t),\end{cases}\end{split}$
for $t=t_{0}+1,t_{0}+2,\cdots$, where $W^{out}:=AX^{-1}$ and $A:=(a_{k(1)},a_{k(2)},\cdots,a_{k(N)})$. Rigorously we regard $\bar{u}(t-L-1)$ as $\bar{y}(t-L-1)$ for $t=t_{0}+1,t_{0}+2\cdots,t_{0}+L$. Note that, since $\bar{r}(t)=Xe_{n_{t^{\prime}}}$ for some $t^{\prime}$ ($t^{\prime}$ is not necessarily equal to $t$), we see $\bar{u}(t)=W^{out}\bar{r}(t)=Ae_{n_{t^{\prime}}}$. Now we show that $\bar{u}_{K}(t)$ tends to a discretized periodic function for any $K$. By (3), for each $K$, we have a chain of the permutation operator $(\sigma_{j},a_{k(j)})$ as follows:
@@
Note that the sequence $j_{1},j_{2},j_{3}\cdots$ is different from $n_{1},n_{2},n_{3}\cdots$ in (4). Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite, then there exist $m,M\in\mathbb{Z}_{\geq 2}$ ($m<M$) such that $(\sigma_{j_{m}},a_{k(j_{m})})=(\sigma_{j_{M}},a_{k(j_{M})})$. This means that there exists $T_{K}$ such that $\bar{u}_{K}(t)$ is a periodic chain for $t>T_{K}$. Moreover $\lim_{K\to\infty}\bar{u}_{K}=:u$ is uniform convergence due to (4). Thus $u(t)$ tends to an almost periodic function for $t\to\infty$ (see Corduneanu [4, (1.2)]).
@@ 4. Appendix
Acknowledgments. The authors are grateful to Professors Masanobu Inubushi and Yoshitaka Saiki for valuable comments. Also the first author thanks J. C. for leading him to join this work. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 20H01819. Research of CN was partly supported by the JSPS Grants-in-Aid for Scientific Research 21K03199.